Find the absolute value of |8+p|=2p-3
Found 2 solutions by ptth176, 100104613:
Answer by ptth176(50)

|8+p|=2p-3
=>8+p=2p-3 or 8+p=-(2p-3)
=>p=11 or p=-2/3

Answer by 100104613(1)

|8+p|=2p-3
=8+p=2p-3
-1 -1
8 =P-3
+8 +8
( 11=p )